# Unitary Designs of Symmetric Local Random Circuits

### v1

@@ Appendix A Application of Theorem 2 to concrete symmetries
In this appendix, we explain how we can apply the result of Theorem 2 to get Theorem 1 in the $\mathbb{Z}_{2}$ and $\mathrm{SU}(2)$ symmetries. First, as for the $\mathbb{Z}_{2}$ symmetry, since $\mathbb{Z}_{2}$ is commutative, every irreducible representation is one-dimensional, and corresponds to the eigenvalue of $\mathrm{Z}^{\otimes n}$. Thus $R$ defined by Eq. (1) is decomposed into two irreducible representations $R_{\lambda}$’s with $\lambda\in\{0,1\}$ defined by $R_{\lambda}(g)=(-1)^{\lambda g}$ for $g\in\{0,1\}$, and the multiplicity $m_{\lambda}$ is given by $2^{n-1}$. We note that every $A\in\mathcal{L}_{k,G,T^{\otimes k}}$ can be written as $A=A_{0}\oplus A_{1}$ with $A_{j}$’s acting on the eigenspaces of $\mathrm{Z}^{\otimes k}$ with eigenvalues $(-1)^{j}$. Since we have $f_{0}(A)=f_{1}(A)=2^{n-k-1}(\mathrm{tr}(A_{0})+\mathrm{tr}(A_{1}))$, we find that $\mathcal{V}$ is spanned by $(1,1)$. Thus Eq. (12) implies that
$t<\min_{\bm{x}\in\{(a,-a)|a\in\mathbb{Z}\backslash\{0\}\}}2^{n-1}(x_{0}^{+}+x_{1}^{+})=2^{n-1}.$ (14)
Next, we consider the $\mathrm{SU}(2)$ case. Since we can apply the method similar to the one in the $\mathrm{U}(1)$ symmetry in the second and the third steps, we only explain the first step of explicitly rewriting Eq. (12). By utilizing the $\mathfrak{sl}(2,\mathbb{C})$ representation theory, the representation $R$ defined by Eq. (3) can be decomposed into the spin-$\lambda$ representations with $\lambda=\{n/2,n/2-1,...,n/2-\lfloor n/2\rfloor\}$. The multiplicity $m_{\lambda}$ is given by $\binom{n}{n/2-\lambda}-\binom{n}{n/2-\lambda-1}$, which corresponds to the difference between the dimensions of the eigenspaces of $\mathrm{Z}^{\mathrm{tot}}$ with eigenvalues $2\lambda$ and $2(\lambda+1)$. In order to take a basis $\{\bm{v}_{j}\}$ of $\mathcal{V}$, we note that $\mathcal{L}_{k,G,T^{\otimes k}}$ can be spanned by permutation operators $Q_{\sigma}$ on $k$ qubits by the Schur-Weyl duality. Thus $\mathcal{V}$ can be spanned by $\bm{f}(Q_{\sigma}\otimes\mathrm{I}^{\otimes n-k})$. By using the property of the binomial coefficients shown in Lemmas 9 and 22 of Ref [42], we can show that $\mathcal{V}$ can be spanned by $\bm{f}([(I-\mathrm{SWAP})/2]^{\otimes j})\otimes\mathrm{I}^{n-2j}$ with $j=0,1,...,\lfloor k/2\rfloor$. Since $[(I-\mathrm{SWAP})/2]^{\otimes j}$ is the projection onto the space of the spin-00 representation on $2j$ spins, $f_{\lambda}([(I-\mathrm{SWAP})/2]^{\otimes j}\otimes\mathrm{I}^{n-2j})$ corresponds to the multiplicity of the spin-$\lambda$ representation in the tensor product of representation of $\mathrm{SU}(2)$ on the rest $n-2j$ qubits, which is given by $\binom{n-2j}{(n-2j)/2-\lambda}-\binom{n-2j}{(n-2j)/2-\lambda-1}$.

### v3

@@ Appendix A Application of Theorem 2 to concrete symmetries
In this appendix, we explain how we can apply the result of Theorem 2 to get Theorem 1 in the $\mathbb{Z}_{2}$ and $\mathrm{SU}(2)$ symmetries. First, as for the $\mathbb{Z}_{2}$ symmetry, since $\mathbb{Z}_{2}$ is commutative, every irreducible representation is one-dimensional, and corresponds to an eigenvalue of $\mathrm{Z}^{\otimes n}$. Thus $R$ defined by Eq. (1) is decomposed into two irreducible representations with $\lambda\in\{0,1\}$ defined by $R_{\lambda}(g)=(-1)^{\lambda g}$ for $g\in\{0,1\}$, and the multiplicity $m_{\lambda}$ is given by $2^{n-1}$. We note that every $A\in\mathcal{L}_{k,G,T^{\otimes k}}$ can be written as $A=A_{0}\oplus A_{1}$ with $A_{j}$’s acting on the eigenspaces of $\mathrm{Z}^{\otimes k}$ with eigenvalues $(-1)^{j}$. Since we have $f_{0}(A)=f_{1}(A)=2^{n-k-1}(\mathrm{tr}(A_{0})+\mathrm{tr}(A_{1}))$, we find that $\mathcal{C}$ is spanned by $(1,1)$. Thus Eq. (13) implies that $t<\min_{\bm{x}\in\{(a,-a)|a\in\mathbb{Z}\backslash\{0\}\}}2^{n-1}(x_{0}^{+}+x_{1}^{+})=2^{n-1}$.
Next, we consider the $\mathrm{SU}(2)$ case. Since we can apply the method similar to the one in the $\mathrm{U}(1)$ symmetry in the second and the third steps, we only explain the first step of explicitly rewriting Eq. (13). By utilizing the $\mathfrak{sl}(2,\mathbb{C})$ representation theory, the representation $R$ defined by Eq. (3) can be decomposed into the spin-$\lambda$ representations with $\lambda=\{n/2,n/2-1,...,n/2-\lfloor n/2\rfloor\}$. The multiplicity $m_{\lambda}$ is given by $\binom{n}{n/2-\lambda}-\binom{n}{n/2-\lambda-1}$, which corresponds to the difference between the dimensions of the eigenspaces of $\mathrm{Z}^{\mathrm{tot}}$ with eigenvalues $2\lambda$ and $2(\lambda+1)$. In order to take a basis $\{\bm{c}_{j}\}$ of $\mathcal{C}$, we note that $\mathcal{L}_{k,G,T^{\otimes k}}$ can be spanned by permutation operators $Q_{\sigma}$ on $k$ qubits by the Schur-Weyl duality. Thus $\mathcal{C}$ can be spanned by $\bm{f}(Q_{\sigma}\otimes\mathrm{I}^{\otimes n-k})$. By using the properties of the binomial coefficients, we can show that $\mathcal{C}$ can be spanned by $\bm{f}([(I-\mathrm{SWAP})/2]^{\otimes j}\otimes\mathrm{I}^{\otimes n-2j})$ with $j=0,1,...,\lfloor k/2\rfloor$ (see Lemma 9 in Ref. [74] for details). Since $[(I-\mathrm{SWAP})/2]^{\otimes j}$ is the projection onto the space of the spin-00 representation on $2j$ spins, $f_{\lambda}([(I-\mathrm{SWAP})/2]^{\otimes j}\otimes\mathrm{I}^{\otimes n-2j})$ corresponds to the multiplicity of the spin-$\lambda$ representation in the tensor product of representation of $\mathrm{SU}(2)$ on the rest $n-2j$ qubits, which is given by $\binom{n-2j}{(n-2j)/2-\lambda}-\binom{n-2j}{(n-2j)/2-\lambda-1}$.